OpenStudy (anonymous):

Solve for x: −2(x + 3) = −2(x + 1) − 4 2 3 All real numbers No solution -------------------------------------------------------------------------------- Question 2 (Multiple Choice Worth 1 points) A social networking site recorded that 80% of the people who became members of the site one month were between 30 and 40 years old. The equation below represents this information, where x represents the total number of people who became members of the site that month. x = 0.8x + 72 How many people who became members of the site that month were between 30 a

OpenStudy (anonymous):

Question 2 (Multiple Choice Worth 1 points) A social networking site recorded that 80% of the people who became members of the site one month were between 30 and 40 years old. The equation below represents this information, where x represents the total number of people who became members of the site that month. x = 0.8x + 72 How many people who became members of the site that month were between 30 and 40 years of age? 112 288 360 432

OpenStudy (kropot72):

Question 1. Expanding both sides gives: -2x - 6 = -2x - 2 - 4 = -2x - 6 Therefore all real numbers is the solution.

OpenStudy (kropot72):

Question 2. x = 0.8x + 72 x - 0.8x = 72 0.2x = 72 x = 72/0.2 = 360 = total number who became members 80% of 360 = 288 = number between 30 and 40 years of age